OpenStudy (anonymous):

What is the length of the hypotenuse in a 45-45-90 triangle that has a leg length of √7?

OpenStudy (anonymous):

a. 7 b. 2√7 c. √21 d. √14

OpenStudy (lucaz):

if you have the 45 45 90 with hypothenuse=1 the legs would be sqrt(2)/2, so you multiply sqrt(2)/2 by x to get the leg of length sqrt(7): sqrt(2)/2 times x = sqrt(7) x = sqrt(14) now to keep te ratio of the triangle you multimply all of the sides by x: 1 times x = 1 times sqrt(14) = sqrt(14)

OpenStudy (lucaz):

they are similar triangles